# Fundamental limitations on the recoverability of quantum processes

### v1

@@ B.4 Proof of Lemma 4
Let $\{e^{Q}_{ij}\}$ denote a canonical basis for the spaces $\mathcal{L}\left(Q\right)$, where $Q=\{A,B,C,D\}$. Then, the above equation leads to
@@ C.1 Proof of monotonicity of relative entropy under quantum superchannel
Let $\mathcal{N}_{A\rightarrow A^{\prime}}$ and $\mathcal{M}_{A\rightarrow A^{\prime}}$ be a quantum channel and a CP map, respectively, and let $\Theta:\mathcal{L}\left(A,A^{\prime}\right)\rightarrow\mathcal{L}\left(C,C^{\prime}\right)$ be a superchannel. To prove that the proposed inequality holds, we start with the left-hand side, i.e., $D\left(\Theta\left(\mathcal{N}\right)\|\Theta\left(\mathcal{M}\right)\right)$. Now, considering a general superchannel $\Theta$, whose action on any CP map is defined in Eq.(13), we can write
@@
$\leq\underset{\Psi_{A^{\prime}EB}}{\mathrm{sup}}D\left(\left(V^{\prime}_{AE\rightarrow C}\hskip 1.42271pt\circ\left(\mathcal{N}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}\;\middle\|\;\left(V^{\prime}_{AE\rightarrow C}\hskip 1.42271pt\circ\left(\mathcal{M}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}\right),$ (101)
where the inequality in the third line follows from the fact that the set of pure states obtained by varying $\Psi_{C^{\prime}B}$ and applying $U_{C^{\prime}\rightarrow A^{\prime}E}$ is smaller in comparison to the set of all pure states in the system $A^{\prime}EB$, i.e, $\{\Psi_{A^{\prime}EB}\}$. In other words, supremum of a set is always greater than or equal to its subsets. Let us define $\left(\left(\mathcal{N}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}:=\rho^{\Psi}_{ABE}$, which is also a density operator since $\mathcal{N}_{A^{\prime}\rightarrow A}$ is a quantum channel, and $\left(\left(\mathcal{M}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}:=\sigma^{\Psi}_{ABE}$, which is a positive operator, because map $\mathcal{M}$ is CP but not trace preserving. Now, we have
$\underset{\Psi_{A^{\prime}EB}}{\mathrm{sup}}D\left(\left(V^{\prime}_{AE\rightarrow C}\hskip 1.42271pt\circ\left(\mathcal{N}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}\;\middle\|\;\left(V^{\prime}_{AE\rightarrow C}\hskip 1.42271pt\circ\left(\mathcal{M}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}\right)$ (102)
$\leq\underset{\Psi_{A^{\prime}EB}}{\mathrm{sup}}D\left(\left(V^{\prime}_{AE\rightarrow C}\otimes\mathrm{id}_{B}\right)\rho^{\Psi}_{ABE}\;\middle\|\;\left(V^{\prime}_{AE\rightarrow C}\otimes\mathrm{id}_{B}\right)\sigma^{\Psi}_{AEB}\right)$ (103)
$\leq\underset{\Psi_{A^{\prime}EB}}{\mathrm{sup}}D\left(\rho^{\Psi}_{ABE}\;\middle\|\;\sigma^{\Psi}_{AEB}\right)$ (104)
$\leq\underset{\Psi_{A^{\prime}EB}}{\mathrm{sup}}D\left(\left(\left(\mathcal{N}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}\;\middle\|\;\left(\left(\mathcal{M}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}\right)=D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right],$ (105)
where, the third line follows from the monotonicity of relative entropy under a quantum channel, and the last line simply follows from the definition of relative entropy between two completely positive maps. Thus, we have the inequality
@@
C.2 Proof of monotonicity of relative entropy under CP and TP preserving supermap
Since $\mathcal{N}$ and $\mathcal{M}$ are quantum channels and the supermap $\Theta$ is CP- and TP-preserving map which implies $\Theta(\mathcal{N})$ and $\Theta(\mathcal{M})$ are also quantum channels. We also have $D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]=D\left({\mathsf{C}}_{\mathcal{N}}^{\Psi}\;\middle\|\;{\mathsf{C}}_{\mathcal{M}}^{\Psi}\right)$ and $D\left[\Theta(\mathcal{N})\;\middle\|\;\mathcal{M}(\Theta)\right]=D\left({\mathsf{C}}_{\Theta(\mathcal{N})}^{\Phi}\;\middle\|\;{\mathsf{C}}_{\Theta(\mathcal{M})}^{\Phi}\right)$, where $\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}$ and $\mathsf{C}_{\Theta\left(\mathcal{M}\right)}^{\Phi}$ are Choi matrices of channels $\Theta\left(\mathcal{N}\right)$ and $\Theta\left(\mathcal{N}\right)$, respectively and $\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}$ and $\mathsf{C}_{\Theta\left(\mathcal{M}\right)}^{\Phi}$ are Choi matrices of channels $\Theta\left(\mathcal{N}\right)$ and $\Theta\left(\mathcal{M}\right)$, respectively (see Appendix D for detail calculations). Since, the Choi operators $\mathsf{C}_{\mathcal{N}}^{\Psi}$, $\mathsf{C}_{\mathcal{M}}^{\Psi}$, $\mathsf{C}_{\Theta(\mathcal{N})}^{\Psi}$ and $\mathsf{C}_{\Theta(\mathcal{N})}^{\Psi}$ are connected via the representing map $\mathfrak{T}$ of the supermap $\Theta$ as $\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}=\mathfrak{T}\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)$ and $\mathsf{C}_{\Theta\left(\mathcal{M}\right)}^{\Phi}=\mathfrak{T}\left(\mathsf{C}_{\mathcal{M}}^{\Psi}\right)$. As the supermap $\Theta$ is CP-preserving, the representing map $\mathfrak{T}$ must be a positive map. However $\mathfrak{T}$ may not be trace-preserving. Replacing it with the following trace preserving map: $\mathfrak{T}^{\prime}(\mathsf{X}):=\mathfrak{T}(\mathsf{X})+\left[\mathrm{tr}(\mathsf{X})-\mathrm{tr}(\mathfrak{T}(\mathsf{X}))\right]\sigma_{0}$, where $\sigma_{0}\in\mathcal{D}(A\otimes A^{\prime})$ is some fixed state and $\mathsf{X}\in\mathcal{L}(C\otimes C^{\prime})$, we have $\mathfrak{T}^{\prime}(\mathsf{C}^{\Psi}_{\mathcal{N}})=\mathfrak{T}(\mathsf{C}^{\Psi}_{\mathcal{N}})$, and similar equality holds for $\mathsf{C}^{\Psi}_{\mathcal{M}}$ as well. Then (MHR, 17, Theorem 1) leads to
@@ C.2 Proof of monotonicity of relative entropy under CP and TP preserving supermap
If we relax the TP-preserving condition on supermap $\Theta$ and take it to be a CP-preserving map, then the representing map $\mathfrak{T}$ will still be a positive map but we can not define a trace-preserving map $\mathfrak{T}^{\prime}$ such that $\mathfrak{T}^{\prime}(\mathsf{C}^{\Psi}_{\mathcal{N}})=\mathfrak{T}(\mathsf{C}^{\Psi}_{\mathcal{N}})$ for all quantum channels $\mathcal{N}$. But we can find a condition for $\mathfrak{T}$ to be a trace-preserving map. The representing map $\mathsf{T}$ is trace-preserving if $\Theta^{*}$ is completely depolarising preserving map ( see Lemma 6). We want to find a similar condition for $\mathfrak{T}$. The map $\mathfrak{T}$ is trace-preserving if and only if its adjoint is unital, i.e., $\mathfrak{T}^{*}\left(\mathbbm{1}\right)=\mathbbm{1}$, which can also be written as $\mathfrak{T}^{*}\left(\mathsf{C}^{\Psi}_{\Theta^{-1}_{\Psi}(\mathcal{R}_{A\rightarrow B})}\right)=\mathsf{C}^{\Phi}_{\Theta^{-1}_{\Phi}(\mathcal{R}_{C\rightarrow D})}$. Using Eq. (32) we obtain the following condition on supermap $\Theta$
$\Theta^{*}\circ\Theta^{*}_{\Psi}\left(\mathcal{R}_{A\rightarrow B}\right)=\Theta^{*}_{\Phi}\left(\mathcal{R}_{C\rightarrow D}\right).$ (109)
@@ C.3 Proof of Proposition 4
where in the second line we have used the fact that supremum of some function over a bigger set is always greater than its supremum over a smaller set, and in the third line we have used in the additivity of quantum relative entropy, i.e., for positive semidefinite operators $\rho_{1},\rho_{2},\sigma_{1}$, and $\sigma_{2}$, we have, $D\left(\rho_{1}\otimes\rho_{2}\|\sigma_{1}\otimes\sigma_{2}\right)=\mathrm{tr}\left(\rho_{2}\right)D\left(\rho_{1}\|\sigma_{1}\right)+\mathrm{tr}\left(\rho_{1}\right)D\left(\rho_{2}\|\sigma_{2}\right)$, and the last line follows simply from the definition of relative entropy between a quantum channel and a CP map.
@@ C.4 Axioms for state and channel entropy
The above-defined entropy of a quantum state satisfies some physically motivated axioms:
@@
    Under the action of random unitary channels, it increases: let $\mathcal{U}(\rho)=\sum_{i}p_{i}U_{i}\rho(U_{i})^{\dagger}$ be a random unitary channel with $U_{i}$’s being unitary operators. Then $S(\mathcal{U}(\rho))\geq S(\rho)$ holds.
@@
    It is maximum at the maximally mixed state and zero at the pure state.
@@ C.5 Proof of additivity of channel entropy
Let $\mathcal{N}_{A_{1}\rightarrow B_{1}}$, and $\mathcal{M}_{A_{2}\rightarrow B_{2}}$ be two quantum channels $\mathcal{R}_{A_{1}\rightarrow B_{1}}$, and $\mathcal{R}_{A_{2}\rightarrow B_{2}}$ be the corresponding completely depolarising map. It can be shown that additivity will hold if the following equality also holds
To show that the above equality holds it is sufficient to prove that the quantity on the left-hand side is less than or equal to the quantity on the right-hand side and then the other way around is also true. If both inequalities are true, then, both sides must be equal. First, we will show the ``"``"`$\geq$"` ` ≥ " part, i.e.,
$D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{M}}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{R}}_{A_{2}\rightarrow B_{2}}\right]\geq D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\right]+D\left[\mathcal{M}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{2}\rightarrow B_{2}}\right].$ (113)
@@
where $R^{\prime}=RA_{1}$ is relabeling of the system $RA_{1}$. Now, using the definition of the completely depolarising map, the operator $\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}_{A_{1}\rightarrow B_{1}}\right)\sigma_{R^{\prime}A_{1}}$ can be simplified to $\sigma_{R^{\prime}}\otimes\mathbbm{1}_{B_{1}}$. Also, let us define $A_{R^{\prime}B_{1}}:=\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}$, then we observe $A_{R^{\prime}}=\mathrm{tr}_{B_{1}}\left(A_{R^{\prime}B_{1}}\right)=\left(\mathrm{id}_{R^{\prime}}\otimes\mathrm{tr_{B_{1}}}\right)\left(\mathrm{id_{R^{\prime}}}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}=\left(\mathrm{id}_{R^{\prime}}\otimes\mathrm{tr_{B_{1}}}\circ\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}=\rho_{R^{\prime}}$. Thus, we can write
$D\left(\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}_{A_{1}\rightarrow B_{1}}\right)\sigma_{R^{\prime}A_{1}}\right)$ $=-S(A_{R^{\prime}B_{1}})-\mathrm{tr}\left(A_{R^{\prime}}\log A_{R^{\prime}}\right)+D\left(A_{R^{\prime}}\;\middle\|\;\sigma_{R^{\prime}}\right),$ (115)
$D\left(\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}\right)$ $=-S\left(A_{R^{\prime}B_{1}}\right)-\mathrm{tr}\left(A_{R^{\prime}}\left(\log\rho_{R^{\prime}}\right)\right).$ (116)

### v2

@@ B.5 Proof of Lemma 4
Let $\{e^{Q}_{ij}\}$ be a canonical basis for the space $\mathcal{L}\left(Q\right)$, where $Q=\{A,B,C,D\}$. Then, the above equation leads to
@@ C.1 Proof of monotonicity of relative entropy under quantum superchannel
Let $\mathcal{N}_{A^{\prime}\rightarrow A}$ and $\mathcal{M}_{A^{\prime}\rightarrow A}$ be a quantum channel and a CP map, respectively, and let $\Theta:\mathcal{L}\left(A^{\prime},A\right)\rightarrow\mathcal{L}\left(C^{\prime},C\right)$ be a superchannel. To prove that the proposed inequality holds, we start with the left-hand side, i.e., $D\left(\Theta\left(\mathcal{N}\right)\|\Theta\left(\mathcal{M}\right)\right)$. Now, considering a general superchannel $\Theta$, whose action on any CP map is defined in Eq.(13), we can write
@@
$\leq\underset{\Psi_{A^{\prime}EB}}{\mathrm{sup}}D\left(\left(V^{\prime}_{AE\rightarrow C}\hskip 1.42271pt\circ\left(\mathcal{N}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}\;\middle\|\;\left(V^{\prime}_{AE\rightarrow C}\hskip 1.42271pt\circ\left(\mathcal{M}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}\right),$
where the inequality in the third line follows from the fact that the set of pure states obtained by varying $\Psi_{C^{\prime}B}$ and applying $U_{C^{\prime}\rightarrow A^{\prime}E}$ is smaller in comparison to the set of all pure states in the system $A^{\prime}EB$, i.e, $\{\Psi_{A^{\prime}EB}\}$. In other words, supremum of a set is always greater than or equal to supremum over its subsets. Let us define $\left(\left(\mathcal{N}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}:=\rho^{\Psi}_{ABE}$, which is also a density operator since $\mathcal{N}_{A^{\prime}\rightarrow A}$ is a quantum channel, and $\left(\left(\mathcal{M}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}:=\sigma^{\Psi}_{ABE}$, which is a positive operator, because map $\mathcal{M}$ is CP but not trace preserving. Now, we have
$\underset{\Psi_{A^{\prime}EB}}{\mathrm{sup}}D\left(\left(V^{\prime}_{AE\rightarrow C}\hskip 1.42271pt\circ\left(\mathcal{N}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}\;\middle\|\;\left(V^{\prime}_{AE\rightarrow C}\hskip 1.42271pt\circ\left(\mathcal{M}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}\right)$
$=\underset{\Psi_{A^{\prime}EB}}{\mathrm{sup}}D\left(\left(V^{\prime}_{AE\rightarrow C}\otimes\mathrm{id}_{B}\right)\rho^{\Psi}_{ABE}\;\middle\|\;\left(V^{\prime}_{AE\rightarrow C}\otimes\mathrm{id}_{B}\right)\sigma^{\Psi}_{AEB}\right)$
$\leq\underset{\Psi_{A^{\prime}EB}}{\mathrm{sup}}D\left(\rho^{\Psi}_{ABE}\;\middle\|\;\sigma^{\Psi}_{AEB}\right)$
$=\underset{\Psi_{A^{\prime}EB}}{\mathrm{sup}}D\left(\left(\left(\mathcal{N}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}\;\middle\|\;\left(\left(\mathcal{M}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}\right)=D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right],$ (114)
where the third line follows from the monotonicity of relative entropy under a quantum channel, and the last line simply follows from the definition of relative entropy between two completely positive maps. Thus, we have the inequality
@@
C.2 Proof of monotonicity of relative entropy under CP and TP preserving supermaps
Since $\mathcal{N}_{A\rightarrow A^{\prime}}$ and $\mathcal{M}_{A\rightarrow A^{\prime}}$ are quantum channels and the supermap $\Theta:\mathcal{L}\left(A^{\prime},A\right)\rightarrow\mathcal{L}\left(C^{\prime},C\right)$ is CP- and TP-preserving map, we have that $\Theta(\mathcal{N})$ and $\Theta(\mathcal{M})$ are also quantum channels. Here the quantum channels $\mathcal{M},\mathcal{N}$ and supermap $\Theta$ are such that the state that realizes the supremum of relative entropy functionals $D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]$ and $D\left[\Theta(\mathcal{N})\;\middle\|\;\Theta(\mathcal{M})\right]$ are element of $\mathsf{FRank}\left(\widetilde{A}\otimes{A}\right)\cap\mathcal{D}\left(\widetilde{A}\otimes{A}\right)$ and $\mathsf{FRank}\left(\widetilde{C}\otimes{C}\right)\cap\mathcal{D}\left(\widetilde{C}\otimes{C}\right)$, respectively. We then have $D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]=D\left({\mathsf{C}}_{\mathcal{N}}^{\Psi}\;\middle\|\;{\mathsf{C}}_{\mathcal{M}}^{\Psi}\right)$ and $D\left[\Theta(\mathcal{N})\;\middle\|\;\mathcal{M}(\Theta)\right]=D\left({\mathsf{C}}_{\Theta(\mathcal{N})}^{\Phi}\;\middle\|\;{\mathsf{C}}_{\Theta(\mathcal{M})}^{\Phi}\right)$, where $\Psi\in\mathsf{FRank}\left(\widetilde{A}\otimes{A}\right)\cap\mathcal{D}\left(\widetilde{A}\otimes{A}\right)$ and $\Phi\in\mathsf{FRank}\left(\widetilde{C}\otimes{C}\right)\cap\mathcal{D}\left(\widetilde{C}\otimes{C}\right)$ (see Appendix D for detailed calculations). Further, $\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}$ and $\mathsf{C}_{\Theta\left(\mathcal{M}\right)}^{\Phi}$ are Choi states of channels $\Theta\left(\mathcal{N}\right)$ and $\Theta\left(\mathcal{N}\right)$, respectively and $\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}$ and $\mathsf{C}_{\Theta\left(\mathcal{M}\right)}^{\Phi}$ are Choi states of channels $\Theta\left(\mathcal{N}\right)$ and $\Theta\left(\mathcal{M}\right)$, respectively. Note that the Choi states $\mathsf{C}_{\mathcal{N}}^{\Psi}$, $\mathsf{C}_{\Theta(\mathcal{N})}^{\Phi}$ and $\mathsf{C}_{\mathcal{M}}^{\Psi}$, $\mathsf{C}_{\Theta(\mathcal{M})}^{\Phi}$ are connected via the representing map $\mathfrak{T}$ of the supermap $\Theta$ as $\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}=\mathfrak{T}\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)$ and $\mathsf{C}_{\Theta\left(\mathcal{M}\right)}^{\Phi}=\mathfrak{T}\left(\mathsf{C}_{\mathcal{M}}^{\Psi}\right)$. As the supermap $\Theta$ is CP-preserving, the representing map $\mathfrak{T}$ must be a positive map. However, $\mathfrak{T}$ may not be a trace preserving map in general. In the case where it is a trace non-increasing, we can always replace it with the following trace preserving map: $\mathfrak{T}^{\prime}(\mathsf{X}):=\mathfrak{T}(\mathsf{X})+\left[\mathrm{tr}(\mathsf{X})-\mathrm{tr}(\mathfrak{T}(\mathsf{X}))\right]\sigma_{0}$, where $\sigma_{0}\in\mathcal{D}(A\otimes A^{\prime})$ is some fixed state and $\mathsf{X}\in\mathcal{L}(C\otimes C^{\prime})$, and we have $\mathfrak{T}^{\prime}(\mathsf{C}^{\Psi}_{\mathcal{N}})=\mathfrak{T}(\mathsf{C}^{\Psi}_{\mathcal{N}})$. A similar equality holds for $\mathsf{C}^{\Psi}_{\mathcal{M}}$ as well. Then (MHR, 17, Theorem 1) leads to
@@ C.2 Proof of monotonicity of relative entropy under CP and TP preserving supermaps
If we relax the TP-preserving condition on supermap $\Theta$ and take it to be a CP-preserving map, then the representing map $\mathfrak{T}$ will be a positive map but we can not define the trace preserving map $\mathfrak{T}^{\prime}$ such that $\mathfrak{T}^{\prime}(\mathsf{C}^{\Psi}_{\mathcal{N}})=\mathfrak{T}(\mathsf{C}^{\Psi}_{\mathcal{N}})$, for all quantum channels $\mathcal{N}$. However, we can find a condition for $\mathfrak{T}$ to be a trace preserving map. The representing map $\mathsf{T}$ is trace preserving if $\Theta^{*}$ is completely depolarising preserving map (see Lemma 6). We want to find a similar condition for $\mathfrak{T}$. The map $\mathfrak{T}$ is trace preserving if and only if its adjoint is unital, i.e., $\mathfrak{T}^{*}\left(\mathbbm{1}\right)=\mathbbm{1}$, which can also be written as $\mathfrak{T}^{*}\left(\mathsf{C}^{\Psi}_{\Theta^{-1}_{\Psi}(\mathcal{R}_{A\rightarrow A^{\prime}})}\right)=\mathsf{C}^{\Phi}_{\Theta^{-1}_{\Phi}(\mathcal{R}_{C\rightarrow C^{\prime}})}$. Using Eq. (32), we obtain that the following condition on supermap $\Theta$ implies that $\mathfrak{T}$ is trace preserving:
$\Theta^{*}\circ\Theta^{*}_{\Psi}\left(\mathcal{R}_{A\rightarrow A^{\prime}}\right)=\Theta^{*}_{\Phi}\left(\mathcal{R}_{C\rightarrow C^{\prime}}\right).$ (118)
@@ C.3 Proof of Proposition 4
where in the second line we have used the fact that supremum of some function over a bigger set is always greater than its supremum over a smaller set, and in the third line we have used the additivity of quantum relative entropy, i.e., for positive semidefinite operators $\rho_{1},\rho_{2},\sigma_{1}$, and $\sigma_{2}$, we have, $D\left(\rho_{1}\otimes\rho_{2}\|\sigma_{1}\otimes\sigma_{2}\right)=\mathrm{tr}\left(\rho_{2}\right)D\left(\rho_{1}\|\sigma_{1}\right)+\mathrm{tr}\left(\rho_{1}\right)D\left(\rho_{2}\|\sigma_{2}\right)$, and the last line follows simply from the definition of relative entropy between a quantum channel and a CP map.
@@ C.4 Axioms for state and channel entropy
The above definition of entropy of a quantum state satisfies some physically motivated axioms:
@@
    Under the action of random unitary channels, it increases: let $\mathcal{U}(\rho)=\sum_{i}p_{i}U_{i}\rho U_{i}^{\dagger}$ be a random unitary channel with $U_{i}$’s being unitary operators. Then $S(\mathcal{U}(\rho))\geq S(\rho)$ holds.
@@
    It is maximum for the maximally mixed state and zero for all pure states.
@@ C.5 Proof of additivity of channel entropy
Let $\mathcal{N}_{A_{1}\rightarrow B_{1}}$ and $\mathcal{M}_{A_{2}\rightarrow B_{2}}$ be two quantum channels, and $\mathcal{R}_{A_{1}\rightarrow B_{1}}$, $\mathcal{R}_{A_{2}\rightarrow B_{2}}$ be the corresponding completely depolarising map. It can be shown that additivity will hold if the following equality also holds
$D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{M}}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{R}}_{A_{2}\rightarrow B_{2}}\right]=D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\right]+D\left[\mathcal{M}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{2}\rightarrow B_{2}}\right].$
To show that the above equality holds, it is sufficient to prove that the quantity on the left-hand side is less than or equal to the quantity on the right-hand side and then the other way around is also true. If both inequalities are true, then, both sides must be equal. First, we will show the ``"``"`$\geq$"` ` ≥ " part, i.e.,
$D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{M}}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{R}}_{A_{2}\rightarrow B_{2}}\right]\geq D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\right]+D\left[\mathcal{M}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{2}\rightarrow B_{2}}\right].$
@@
where $R^{\prime}=RB_{2}$ is relabeling of the system $RB_{2}$. Now, using the definition of the completely depolarising map, the operator $\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}_{A_{1}\rightarrow B_{1}}\right)\sigma_{R^{\prime}A_{1}}$ can be simplified to $\sigma_{R^{\prime}}\otimes\mathbbm{1}_{B_{1}}$. Also, let us define $A_{R^{\prime}B_{1}}:=\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}$, then we observe $A_{R^{\prime}}=\mathrm{tr}_{B_{1}}\left(A_{R^{\prime}B_{1}}\right)=\left(\mathrm{id}_{R^{\prime}}\otimes\mathrm{tr_{B_{1}}}\right)\left(\mathrm{id_{R^{\prime}}}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}=\left(\mathrm{id}_{R^{\prime}}\otimes\mathrm{tr_{B_{1}}}\circ\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}=\rho_{R^{\prime}}$. Thus, we can write
$D\left(\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}_{A_{1}\rightarrow B_{1}}\right)\sigma_{R^{\prime}A_{1}}\right)$ $=-S(A_{R^{\prime}B_{1}})-\mathrm{tr}\left(A_{R^{\prime}}\log\sigma_{R^{\prime}}\right),$ (122)
$D\left(\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}\right)$ $=-S\left(A_{R^{\prime}B_{1}}\right)-\mathrm{tr}\left(A_{R^{\prime}}\log\rho_{R^{\prime}}\right).$ (123)